Solving a linear equation: Problem type 2
Solve for x. Write your answer as a fraction in simplest form.
4/3x + 3/2 = 7/9
Please help me understand how to do this problem.

Answer by stanbon(48558)

Solve for x. Write your answer as a fraction in simplest form.
4/3x + 3/2 = 7/9
Multiply thru by 18x to get:
6(4)+3(9x)=7(2x)
24+27x=14x
13x=-24
x=-24/13
Cheers,
Stan H.